OpenStudy (anonymous):

solve for x: \[\sqrt{xsquared+8x}= 4\sqrt{3}\] A. -12 B. -4 C. 4 D. -12, 4

OpenStudy (anonymous):

\[\sqrt{x^2+8x}=4\sqrt3\] Square both sides: \[\left(\sqrt{x^2+8x}\right)^2=\left(4\sqrt3\right)^2\\ x^2+8x=48\\ x^2+8x-48=0\] Hint: left side can be easily factored.

OpenStudy (anonymous):

so what would be the answer?

jhonyy9 (jhonyy9):

@SithsAndGiggles so you dont like ending it ?

jhonyy9 (jhonyy9):

Jessica can you factorizing ?

OpenStudy (anonymous):

@jhonyy9, I believe the second principle should answer your question. http://openstudy.com/code-of-conduct